OpenStudy (anonymous):

Is the point for this equation 3,4? 2x + 2y = 14 3x -y = 4

OpenStudy (whpalmer4):

Plug in the numbers and do the arithmetic: \[2(3)+2(4) = 14\]\[6+8=14\checkmark\] but you have to test both equations! \[3(3)-(4) = 4\]\[9-4\ne4\]Oops. Not a solution of the two equations.

OpenStudy (anonymous):

so what are the points?

OpenStudy (whpalmer4):

\[2x+2y=14\]\[3x-y=4\]Multiply the second equation by 2 and add the two equations together: \[2x+2y=14\]\[6x-2y=8\]-------------- \[8x+0y=22\] Solve that for \(x\), then determine \(y\) with either of the original equations. I'll check your results for you.